OpenStudy (anonymous):

Solve using the quadratic formula. Approximate answers to the nearest tenth. 2x^2-3x-8=0

OpenStudy (austinl):

We are given the equation, \(ax^2+bx+c=0\) The quadratic formula is, \(x=\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\) Think you could solve from there?